OpenStudy (anonymous):

(Y-3)/(x-4)=10 solve for y

OpenStudy (hawkfalcon):

\[\frac{ (y-3) }{ x-4} = 10\] \[y-3 = 10(x-4)\]

OpenStudy (hawkfalcon):

\[y-3 = 10x -40\]